# Numerical investigation of buoyancy-aided mixed convective flow past a square cylinder inclined at 45 degrees

### v1

@@ Abstract
The present study numerically investigates two-dimensional mixed convective flow of air past a square cylinder placed at an angle of incidence of $\alpha=45^{\circ}$ to the free-stream. We perform direct numerical simulations (DNS) for a Reynolds number (Re) of 100 and a range of Richardson numbers (Ri) between 0.0 and 1.0 and a Prandtl number (Pr) of 0.7. The critical Richardson number at which the near-field becomes a steady flow from an unsteady one, using Stuart-Landau analysis, is found to be Ri $=0.68$, and simultaneously, the far-field unsteadiness emerges. There is no range of Ri for which the entire flow field is seen to be steady. At a relatively moderate Ri, the flow field reveals the presence of vorticity inversion through the momentum deficit/addition in the downstream region. We discuss the dual wake-plume nature of the flow beyond the cylinder. The wake exhibits characteristics similar to those of a buoyant jet in the far-field at increased buoyancy. We explore the cause of the far-field unsteadiness, and discuss the mechanism of the observed flow physics using instantaneous and time-averaged flow fields. The important flow quantities, such as force coefficients, vortex shedding frequency, and Nusselt number, are discussed at various Richardson numbers.
@@ I Introduction
Flow past bluff bodies is of significant academic and practical interest. Offshore structures, aircraft flows, tall buildings, and heat exchangers are a few situations where understanding flow past obstacles is crucial. Many phenomena, such as Bénard-von Kármán vortex street formation, vortex shedding, oscillation of the shear layer, and its eventual detachment, have been observed in such flows. The wake dynamics are complex due to their dependence on multiple parameters, including the aspect ratio, blockage ratio, and boundary conditions. Such flows demonstrate the traits of nonlinear dynamical systems because the unsteady Navier-Stokes equations are highly nonlinear, even at low Reynolds numbers (Re).
@@
It is now a well known fact that bluff body flows reveal a series of transitions. At Re $\approx$ 5, the flow past a square cylinder at angle of incidence $\alpha=0^{\circ}$ separates from the downstream corners and reattaches after a finite length in the wake, whose length increases with an increase in Re. The flow undergoes a Hopf bifurcation (steady to unsteady flow) at Re $\approx$ 45. [1, 2] In the case of $\alpha=45^{\circ}$, the Hopf bifurcation occurs at Re $=42$, due to the lower bluffness compared to the $\alpha=0^{\circ}$ case. [3] For Re in the range (150-175), a second (spatial) transition takes the two-dimensional flow towards three-dimensionality. Further increase in Re takes the flow to turbulence along the Ruelle-Takens-Newhouse route to chaos, as shown by Robichaux et al. [4, 5] All computations in this study are two-dimensional, as the transition to three-dimensionality occurs well beyond Re $=150$, while our study is for Re $=100$. Remarkably, despite geometrical differences, the wake characteristics of square and circular cylinders exhibit striking similarities. [6, 7, 8, 9]
@@
In the computational modeling of any free shear flow, a significant challenge is resolving a physically infinite domain into a computationally finite one. Due to computational limitations, it is not possible to use an infinite domain. As a result, we must impose artificial boundaries, which raises an issue with the domain ”finiteness,” quantified by blockage ratio ($\beta$). The blockage ratio, defined as $\beta$ = $d/H$, is the ratio of the projected length of the cylinder ($d$) to the transverse length of the domain ($H$). Sohankar et al. [3] reported that reduction of blockage ratio causes drag coefficient ($\overline{C}_{D}$), pressure drag coefficient ($\overline{C}_{P}$), and Strouhal number (St) to decrease. Experimental studies have been conducted by many authors [10, 11, 12] and the effect of flow confinement has been studied in detail. [13, 14, 15, 16]
@@
The heating of the obstacle in such flows leads to the emergence of numerous new phenomena. In mixed convective flows, the heated cylinder generates density differences in the fluid, adding momentum to the near-field that has a momentum deficit, which ”suppresses” vortex shedding. [17, 18, 19] Sharma and Eswaran report that the flow field is steady beyond a critical Richardson number (Ri) of 0.15. [20, 17, 21] They have also studied the effect of buoyancy-aided (Ri $>0.0$) and buoyancy-opposed (Ri $<0.0$) configurations, where the cylinder is at a higher or lower temperature than the free-stream, respectively. Studies have also investigated flow in the regimes of forced and natural convection [22, 23] as well as natural convection flows with other geometries, which reveal a similar route to chaos as that of forced flows past cylinders. [24, 25, 26] The effect of angle of incidence has been studied for isothermal [3, 27] and non-isothermal flows. [28, 29] There is an effect of the downstream length of the domain on the flow physics, as shown by Dushe, [17] who reported that the increased strength of the buoyancy far downstream causes previously unexamined physical phenomena to take place. Most studies have not considered a domain longer than $40d$ (d is the projected length of the cylinder), leading to the loss of far-field flow physics. Motivated by the work of Dushe, we have considered an exceptionally long domain of $110d$ downstream of the cylinder. For ease of further discussion, we divide the domain beyond the cylinder into three distinct regions: (i) near-field (Y $<15$), (ii) intermediate-field ($15\leq$ Y $\leq 60$) and (iii) far-field (Y $>60$).
@@
Most studies on mixed convective flows past bluff bodies have dealt with the near-field. [30, 28, 29, 21] There has been no numerical investigation on the far-field characteristics of mixed convective flows past cylinders except that of Dushe, [17] who has performed a detailed far-field investigation for the $\alpha=0^{\circ}$ case and found further transitions in the far-field at higher Ri. Only one study exists for the $\alpha=45^{\circ}$ case in the buoyancy-aided regime, by Arif and Hasan, [28] who have discussed only the near-field flow. An experimental study by Kimura and Bejan [31] provides qualitative evidence of the presence of unsteadiness in the far-field of natural convection plumes. For $\alpha=45^{\circ}$, no work exists that discusses the physics of the near-field and far-field flow in the buoyancy-aided flow regime, motivating us to perform this study to observe the far-field characteristics as well as examine the effect of the cylinder orientation on the nature of the downstream flow.
@@
In forced flows, the far-field becomes steady due to the diffusion of the shed vortices. Since we are considering buoyancy-aided flow, it may so happen that the natural convective part of the flow may dominate over the forced convective part in the far-field. The reduced strength of the wake, combined with the increased strength of the plume arising from natural convection, may lead to large-scale unsteadiness in the far-field, which is supported by our results and those of Dushe. [17] Therefore, our objective is to investigate the mixed convective flow past an inclined square cylinder at $\alpha=45^{\circ}$, with the aim of investigating the near-field and far-field dynamics in detail.
@@
To achieve the objectives outlined above, we conduct direct numerical simulations (DNS) of two-dimensional unconfined, upward, mixed convective flow of air past a square cylinder (heated to a constant wall temperature higher than the free-stream temperature) at a fixed incidence angle $\alpha=45^{\circ}$ to the free-stream. The present investigation considers flow in the buoyancy-aided regime as the buoyancy force assists the flow in the upward direction. The flow parameters used are Re $=100$ and $0.0\leq Ri\leq 1.0$. Air has been considered as the fluid, and the Prandtl number (Pr) is taken as 0.7.
@@
The remaining manuscript is structured as follows: In Section II, we discuss the numerical method, governing equations and boundary conditions, details of the validation, grid independence test, and domain independence test. In Section III, we present the results of our investigation which include integral parameters, instantaneous, and time-averaged results. These results are followed by a discussion on the inversion of vorticity, suppression of vortex shedding, and the far-field unsteadiness. Finally, the main findings are summarized in Section IV.
@@ II.2 Governing differential equations and boundary conditions
$Gr$ is the Grashof number, $g$ is the acceleration due to gravity, $\beta$ is the coefficient of volumetric thermal expansion, $Pr$ is the Prandtl number of the fluid, $\nu$ is the kinematic viscosity of the fluid, $\rho$ is the density of the fluid, $\mu$ is the dynamic viscosity of the fluid, $C_{ph}$ is the specific heat of the fluid, and $K$ is the thermal conductivity of the fluid.
Air is the fluid under consideration and the Prandtl number is taken as 0.7. The effect of viscous dissipation has been neglected. The Boussinesq approximation has been utilized to convert density differences to corresponding temperature differences, and the buoyancy force term has been included in the $y$-momentum equation. In Equations 1, 2 and 3, all velocities and lengths are non-dimensionalized with the free-stream velocity $V_{\infty}$ and the projected length of the cylinder in the streamwise direction $d$, respectively. Similarly, time and pressure are non-dimensionalized with $d/V_{\infty}$ and $\rho V_{\infty}^{2}$, respectively. Temperature is non-dimensionalized as follows:
@@
Transverse confining surfaces ($X=\pm H/2$) are unbounded and subsequently, they are modeled as free-slip, adiabatic surfaces.
@@
At the outlet ($Y=L$), the selection of an appropriate outflow boundary condition is challenging because there is no unique prescription for the outflow boundary of external flows. It has been shown that the selection of the outflow boundary condition greatly affects upstream flow in external flows. Therefore, it should be chosen such that there is minimal effect on upstream flow. Accordingly, a convective boundary condition, first prescribed by Orlanski, [32] has been employed for both velocity and temperature:
@@
(b) Near the cylinder vicinity ($-1\leq$ X $\leq 1$, $-1\leq$ Y $\leq 1$)
@@ II.3 Solution Methodology
The numerical method used in the present work is an improved version of the finite-difference Marker and Cell algorithm proposed by Harlow and Welch. [33] The governing equations have been solved on a staggered grid without encountering the decoupling of pressure and velocity. Velocity components are located on the cell face to which they are normal and pressure is located at the cell center. Velocity components are interpolated to the cell center when required. Convective and diffusive terms are discretized using a second-order central difference scheme. Time advancement of both convective and diffusive terms is done using a second-order time accurate explicit Adams-Bashforth method.
@@
The corrector step (Equation 7) is solved using point Gauss-Seidel method with an over relaxation parameter $\omega_{0}$, which accelerates pressure correction. An optimum value of $\omega_{0}$ = 1.8 has been used for all calculations in the present work.
The grid utilized is shown in Figure 2. It is a non-uniform, structured, Cartesian grid with a fine meshes in the near-wall region. In the vicinity of the cylinder, the grid is uniform with a cell size of $\Delta x=\Delta y=0.01$. Beyond a certain distance from the obstacle in all directions, the grid spacing is increased in a proper ratio to minimize computational cost.
@@
Nusselt number (Nu) can be written as:
@@
Table 2: Comparison of mean Nusselt number at Re $=100$ ($\alpha=0^{\circ}$).
@@
Present Sharma and Eswaran [20]
0.0 3.87 4.09
0.25 4.12 4.30
0.50 4.37 4.55
0.75 4.55 4.73
1.0 4.70 4.87
@@
The accuracy of the computed solution depends on the discretization error which is controlled by the grid resolution. To ensure that the grid is sufficiently fine to reduce the discretization error, the grid independence test has been carried out using three different grid sizes namely, G1 (619 $\times$ 320), G2 (777 $\times$ 396), and G3 (883 $\times$ 504) at Re $=100$ and Ri $=1.0$ by comparing the same parameters used in the validation. The parameters computed using the three grid sizes along with the minimum grid spacing in the vicinity of the cylinder are presented in Table 5. The domain size is kept constant for the grid independence test (120 $\times$ 50). Critical Ri has also been found for all three grids. In G1, it is 0.74, and in G2 and G3, it is 0.68. We observe that G2 produces results with minimal deviation from G3. The mean drag coefficient reveals an error of $2.82\%$ and $1.23\%$ for G1 and G2, respectively, with respect to G3. The corresponding error in case of Strouhal number is $2.09\%$ and $-5.88\%$, and for mean Nusselt number, it is $1.09\%$ and $0.55\%$, respectively. In view of moderate computational cost and closeness of results with G3, the grid G2 is utilized for all subsequent calculations.
@@
As mentioned in Section I, integral parameters depend on the blockage ratio. To ensure that the calculated integral parameters are free from dependence on blockage ratio, the domain independence test has been carried out using two different domains, namely D1 ($120\times 50$, $\beta=2\%$) and D2 ($120\times 67$, $\beta=1.5\%$) at Re $=100$ and Ri $=1.0$ by comparing the same parameters that are used in the validation. Both domains have the same streamwise extent. As the original domain is already very long, it is expected that an increase in its length downstream of the cylinder will not change the flow physics. A detailed investigation on the effect of the domain length has been conducted by Dushe. [17] The parameters calculated using the two domains are presented in Table 6. The mean drag coefficient reveals an error of $0.4\%$ for D1 with respect to D2, whereas for mean Nusselt number and Strouhal number, it is $0.6\%$, and $6.8\%$, respectively. The domain independence test reveals that the results do not depend on blockage. Therefore, domain D1 is used for all subsequent calculations.
Table 6: Mean drag coefficient, mean Nusselt number and Strouhal number for different domain sizes at Re $=100$ and Ri $=1.0$.
Domain Mean drag coefficient ($\overline{C}_{D}$) Mean Nusselt number ($\overline{Nu}$) Strouhal number (St)
D1 ($120\times 50$) 2.44 5.47 0.221
D2 ($120\times 67$) 2.45 5.44 0.237

### v3

@@ Abstract
The present study numerically investigates the two-dimensional mixed convective flow of air past a square cylinder placed at an angle of incidence, $\alpha=45^{\circ}$ to the free-stream. We perform direct numerical simulations (DNS) for a Reynolds number (Re) of 100, a range of Richardson numbers (Ri) between 0.0 and 1.0, and a Prandtl number (Pr) of 0.7. The critical Richardson number at which the near-field becomes a steady flow from an unsteady one lies between $0.65$ and $0.7$, along with a simultaneous emergence of the far-field unsteadiness. There is no range of Ri for which the entire flow field is seen to be steady. At a relatively moderate Ri, the flow field reveals the presence of vorticity inversion through the momentum addition in the downstream region. We discuss the dual wake-plume nature of the flow beyond the cylinder. The wake exhibits characteristics similar to those of a buoyant jet in the far-field at increased buoyancy. We explore the cause of the far-field unsteadiness and discuss the mechanism of the observed flow physics using instantaneous and time-averaged flow fields. The important flow quantities, such as force coefficients, vortex shedding frequency, and Nusselt number, are discussed at various Richardson numbers.
@@ I Introduction
Flow past bluff bodies is of significant academic and practical interest. Offshore structures, aircraft flows, tall buildings, and heat exchangers are a few situations where understanding of the flow past obstacles is crucial. Many phenomena, such as Bénard-von Kármán vortex street formation, oscillation of the shear layer, and its eventual detachment, have been observed in such flows. The wake dynamics are complex due to their dependence on multiple parameters, including the aspect ratio, blockage ratio, and boundary conditions. Such flows demonstrate the traits of nonlinear dynamical systems because the unsteady Navier-Stokes equations are highly nonlinear, even at low Reynolds numbers (Re).
@@
It is now well-known that bluff body flows reveal a series of transitions. At Re $\approx$ 5, the flow past a square cylinder at an angle of incidence $\alpha=0^{\circ}$ separates at the front sharp edges, and the two separating shear layers merge downstream of the cylinder forming a symmetric steady recirculation region, whose length increases with an increase in Re. The flow undergoes a Hopf bifurcation (steady to unsteady flow) at Re $\approx$ 45 [1, 2]. Depending on Re, these two shear layers sometimes reattach on the side walls before finally separating at the rear edges of the cylinder. For Re in the range (150-175), a second (spatial) transition takes the two-dimensional flow towards three-dimensionality. Further, an increase in Re takes the flow to turbulence along the Ruelle-Takens-Newhouse route to chaos, as shown by Robichaux et al. [3, 4]. All computations in this study are two-dimensional, as the transition to three-dimensionality occurs well beyond Re $=150$, while our study is for Re $=100$. Remarkably, despite geometrical differences, the wake characteristics of square and circular cylinders exhibit striking similarities [5, 6, 7, 8].
In the case of $\alpha=45^{\circ}$, the Hopf bifurcation occurs at Re $=42$ due to the lower bluffness compared to the $\alpha=0^{\circ}$ case [9]. This is the lowest critical Reynolds number among all angles of incidence in the range $0^{\circ}<\alpha<45^{\circ}$. The second transition occurs around Re $\approx$ 130 as reported by Yoon et al. [10] where there is an onset of three-dimensionality in the flow. Once again, this critical value is the lowest among all angles of incidence in the range $0^{\circ}<\alpha<45^{\circ}$. While the sequence of flow transitions remains identical between $0^{\circ}$ and $45^{\circ}$, the critical Re for each transition is significantly lower in the $\alpha=45^{\circ}$ case.
@@
In the computational modeling of any free shear flow, a significant challenge is resolving a physically infinite domain into a computationally finite one. Due to computational limitations, it is not possible to use an infinite domain. As a result, we must impose artificial boundaries, which raises an issue with the domain ”finiteness,” quantified by blockage ratio ($\beta$). The blockage ratio, defined as $\beta$ = $d/H$, is the ratio of the projected length of the cylinder ($d$) to the transverse length of the domain ($H$). Sohankar et al. [9] reported that a reduction in blockage ratio causes drag coefficient ($\overline{C}_{D}$), pressure drag coefficient ($\overline{C}_{P}$), and Strouhal number (St) to decrease. Many authors have conducted experimental studies [11, 12, 13], and the effect of flow confinement has been studied in detail [14, 15, 16, 17].
@@
The heating of the obstacle in such flows leads to numerous new phenomena. In mixed convective flows, the heated cylinder generates density differences in the fluid, adding momentum to the near-field that would otherwise have a momentum deficit due to the presence of the cylinder, which ”suppresses” vortex shedding [18, 19, 20]. Sharma and Eswaran reported that the flow field is steady beyond a critical Richardson number (Ri) of 0.15 [21, 18, 22]. They have also studied the effect of buoyancy-aided (Ri $>0.0$) and buoyancy-opposed (Ri $<0.0$) configurations, where the cylinder is at a higher or lower temperature than the free-stream, respectively. Studies have also investigated flow in the regimes of forced and natural convection [23, 24] as well as natural convection flows with other geometries, which reveal a similar route to chaos as that of forced flows past cylinders [25, 26, 27]. The effect of angle of incidence has been studied for isothermal [9, 10] and non-isothermal flows [28, 29]. There is an effect of the downstream length of the domain on the flow physics, as shown by Dushe [18], who reported that the increased strength of the buoyancy far downstream causes previously unexamined physical phenomena to take place. Most studies have not considered a domain longer than $50d$ ($d$ is the projected length of the cylinder), leading to the loss of far-field flow physics. Motivated by the work of Dushe, we have considered an exceptionally long domain of $110d$ downstream of the cylinder. For ease of further discussion, we have divided the domain beyond the cylinder into three distinct regions: (i) near-field (Y $<15$), (ii) intermediate-field ($15\leq$ Y $\leq 60$), and (iii) far-field (Y $>60$).
@@
Most studies on mixed convective flows past bluff bodies have dealt with the near-field [30, 28, 29, 22]. There has been no numerical investigation on the far-field characteristics of mixed convective flows past cylinders except that of Dushe [18], who has performed a detailed far-field investigation for the $\alpha=0^{\circ}$ case and found further transitions in the far-field at higher Ri. Only one study exists for the $\alpha=45^{\circ}$ case in the buoyancy-aided regime, by Arif and Hasan, [28] who have discussed only the near-field flow. An experimental study by Kimura and Bejan [31] provides qualitative evidence of unsteadiness in the far-field of a natural convection plume resulting from a point heat source. The heat is increased progressively, and changes in the plume structure in the far-field are qualitatively examined. For $\alpha=45^{\circ}$, no work exists that discusses the physics of both the near-field and far-field flow in the buoyancy-aided flow regime and, thus, is the motivation to perform this study to observe the far-field characteristics as well as examine the effect of the cylinder orientation on the nature of the downstream flow.
@@
In forced flows, the far-field becomes steady due to the diffusion of the shed vortices. Since we are considering buoyancy-aided flow, the natural convective part of the flow may dominate over the forced convective part in the far-field. The reduced strength of the wake, combined with the increased strength of the plume arising from natural convection, may lead to unsteadiness in the far-field, which is supported by our results and those of Dushe [18]. Therefore, our objective is to investigate the mixed convective flow past an inclined square cylinder at $\alpha=45^{\circ}$, with the aim of investigating the near-field and far-field dynamics in detail.
@@
To achieve the objectives outlined above, we have conducted direct numerical simulations (DNS) of two-dimensional, unconfined, buoyancy-aided (vertically upward flow with gravity acting downward) mixed convective flow of air past a heated square cylinder (heated to a constant wall temperature higher than the free-stream temperature) at a fixed incidence angle, $\alpha=45^{\circ}$ to the free-stream. The present investigation considers flow in the buoyancy-aided regime as the buoyancy force assists the flow in the upward direction. The flow parameters used are Re $=100$ and $0.0\leq Ri\leq 1.0$. Air has been considered as the fluid, the Prandtl number (Pr) of which is taken as 0.7.
@@
The remaining manuscript is structured as follows: In Section II, we have discussed the numerical method, governing equations and boundary conditions, details of the validation, grid independence test, and domain independence test. In Section III, we have presented the results of our investigation, which include integral parameters, instantaneous, and time-averaged results. These results are followed by a discussion on the inversion of vorticity, suppression of vortex shedding, and far-field unsteadiness. Finally, the main findings are summarized in Section IV.
@@ II.2 Governing Differential Equations and Boundary Conditions
$Gr$ is the Grashof number, $g$ is the acceleration due to gravity, $\beta$ is the coefficient of volumetric thermal expansion, $Pr$ is the Prandtl number of the fluid, $\nu$ is the kinematic viscosity of the fluid, $\rho$ is the density of the fluid, $\mu$ is the dynamic viscosity of the fluid, $C_{ph}$ is the specific heat of the fluid, and $K$ is the thermal conductivity of the fluid. $\delta_{i,j}$ is the Kronecker delta.
Air is the fluid under consideration, and the Prandtl number is taken as 0.7. The effect of viscous dissipation has been neglected. The Boussinesq approximation has been utilized to convert density differences to corresponding temperature differences, and the buoyancy force term has been included in the $y$-momentum equation. In Equations 1, 2 and 3, all velocities and lengths are non-dimensionalized with the free-stream velocity $V_{\infty}$ and the projected length of the cylinder in the streamwise direction $d$, respectively. Similarly, time and pressure are non-dimensionalized with $d/V_{\infty}$ and $\rho V_{\infty}^{2}$, respectively. Temperature is non-dimensionalized as follows:
@@
Transverse confining surfaces ($X=\pm H/2$) are unbounded, and subsequently, they are modeled as free-slip, adiabatic surfaces.
@@
At the outlet ($Y=L$), selecting an appropriate outflow boundary condition is challenging because there is no unique prescription for the outflow boundary of external flows. It has been shown that the selection of the outflow boundary condition greatly affects the upstream flow, particularly in subsonic flow. Therefore, it should be chosen to have minimal effect on upstream flow. Accordingly, a convective boundary condition, first prescribed by Orlanski [32], has been employed for both velocity and temperature:
@@
(b) Near the cylinder vicinity ($-0.85\leq$ X $\leq 0.85$, $-1\leq$ Y $\leq 1$)
@@ II.3 Solution Methodology
The numerical method used in the present work is an improved version of the finite-difference Marker and Cell algorithm proposed by Harlow and Welch [33]. The governing equations have been solved on a staggered grid that helps in the coupling of pressure and velocity fields. Velocity components are located on the cell face to which they are normal and pressure is located at the cell center. Velocity components are interpolated to the cell center when required. Convective and diffusive terms are discretized using a second-order central difference scheme. Time advancement of both convective and diffusive terms is done using a second-order time-accurate explicit Adams-Bashforth method.
@@
The corrector step (Equation 7) is solved using the point Gauss-Seidel method with an over-relaxation parameter $\omega_{0}$, which accelerates pressure correction. An optimum value of $\omega_{0}$ = 1.8 has been used for all calculations in the present work.
The grid utilized is shown in Figure 2. It is a non-uniform, structured Cartesian grid with fine meshes in the near-wall region. In the vicinity of the cylinder, the grid is uniform with a cell size of $\Delta x=\Delta y=0.01$. Beyond a certain distance from the obstacle in all directions, the grid spacing is increased in a proper ratio to minimize computational cost.
@@
The heat transfer balance at the solid-liquid interface is given by
@@
Using the above relation, Nusselt number (Nu) can be written as:
@@
Table 2: Comparison of mean Nusselt number and drag coefficient at Re $=100$ ($\alpha=0^{\circ}$).
@@
Mean drag coefficient
($\overline{C}_{D}$)
Present Sharma and Eswaran [21] Present Sharma and Eswaran [21]
0.0 3.90 4.09 1.48 1.57
0.25 4.12 4.30 1.66 1.76
0.50 4.37 4.55 1.98 2.07
0.75 4.55 4.73 2.28 2.36
1.0 4.70 4.87 2.57 2.63
@@
The accuracy of the computed solution depends on the discretization error, which is controlled by the grid resolution. To ensure that the grid is sufficiently fine to reduce the discretization error, a grid independence test has been carried out using three different grid sizes, namely, G1 (619 $\times$ 320), G2 (777 $\times$ 396), and G3 (883 $\times$ 504) at Re $=100$ and Ri $=1.0$ by comparing the same parameters used in the validation. The parameters computed using the three grid sizes, along with the minimum grid spacing in the vicinity of the cylinder, are presented in Table 5. The domain size is kept constant for the grid independence test (120 $\times$ 50). We observe that G2 produces results with minimal deviation from G3. The mean drag coefficient reveals an error of $2.82\%$ and $1.23\%$ for G1 and G2, respectively, with respect to G3. The corresponding error in the case of the Strouhal number is $2.09\%$ and $5.88\%$, and for the mean Nusselt number, it is $1.09\%$ and $0.55\%$, respectively. Given the moderate computational cost and closeness of results with G3, the grid G2 is utilized for all subsequent calculations.
@@
As mentioned in Section I, integral parameters depend on the blockage ratio. To ensure that the calculated integral parameters are free from dependence on the blockage ratio, a domain independence test has been carried out using two different domains, namely D1 ($120\times 50$, $\beta=2\%$) and D2 ($120\times 67$, $\beta=1.5\%$) at Re $=100$ and Ri $=1.0$ by comparing the same parameters that are used in the validation. Both domains have the same streamwise extent. As the original domain is already very long, it is expected that an increase in its length downstream of the cylinder will not change the flow physics. A detailed investigation on the effect of the domain length has been conducted by Dushe [18]. The mean drag coefficient ($\overline{C}_{D}$), mean Nusselt number ($\overline{Nu}$), and Strouhal number (St) for the domain D1 are 2.44, 5.47, and 0.221. The corresponding values for the domain D2 are 2.45, 5.44, and 0.237. The mean drag coefficient reveals an error of $0.4\%$ for D1 with respect to D2, whereas for mean Nusselt number and Strouhal number, it is $0.6\%$, and $6.8\%$, respectively. The domain independence test reveals that the results do not depend on blockage. Therefore, domain D1 is used for all subsequent calculations.